OpenStudy (anonymous):

4. Evaluate the expression for a = –1 and b = 5. (ab)2

OpenStudy (anonymous):

@texaschic101 help with this one

OpenStudy (agent47):

(-1*5)^2 = (-5)^2=25

OpenStudy (anonymous):

the ab is confusing me

OpenStudy (anonymous):

multiply -1 and 5

OpenStudy (texaschic101):

agent47 is correct

OpenStudy (anonymous):

ok Break it down for me

undeadknight26 (undeadknight26):

yes its really just (a x b) x 2 unless the actually equation is (a x b)^2...

OpenStudy (texaschic101):

oh....I assumed it was squared...is it ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-1 * 5= -5 * 5 = 25

OpenStudy (anonymous):

ab^2

OpenStudy (anonymous):

-5* sorry

OpenStudy (texaschic101):

ab is the same as a x b -1 * 5 = -5 and -5^2 = (-5)(-5) = 25

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

yeah I got it now.

OpenStudy (texaschic101):

good to hear :)